# Quantum-gas microscopy of the Bose-glass phase

### v1

@@ Abstract
Disordered potentials fundamentally alter the transport properties and coherence of quantum systems. They give rise to phenomena such as Anderson localization in non-interacting systems, inhibiting transport. When interactions are introduced, the interplay with disorder becomes significantly more complex, and the conditions under which localization can be observed remain an open question. In interacting bosonic systems, a Bose glass is expected to emerge at low energies as an insulating yet compressible state without long-range phase coherence. While originally predicted to occur as a ground-state phase, more recent studies indicate that it exists at finite temperature. A key open challenge has been the direct observation of reduced phase coherence in the Bose-glass regime. In this study, we utilize ultracold bosonic atoms in a quantum-gas microscope to probe the emergence of the Bose-glass phase in a two-dimensional square lattice with a site-resolved, reproducible disordered potential. We identify the phase through in-situ distribution and particle fluctuations, via a local measurement of the Edwards-Anderson parameter. To measure the short-range phase coherence in the Bose glass, we employ Talbot interferometry in combination with single-atom-resolved detection. Finally, by driving the system in and out of the Bose-glass phase, we observe signatures for non-ergodic behavior.
@@
Disordered potentials govern electron transport and thermalization in real materials [1, 2, 3], and have also been studied in many-body quantum systems with ultracold atoms [4, 5, 6, 7], in both experiment and theory. For non-interacting quantum systems, destructive interference from scattering in the disordered medium leads to Anderson localization [8], a phenomenon that has been observed for example in photonic lattices [9, 10] and in cold-atom systems [11, 12, 13, 14]. In general, the addition of interactions gives rise to a complex set of phenomena including, at low energies, the emergence of the Bose glass [15, 16, 17, 18] and spin glasses [19] in bosonic and fermionic systems. The Bose glass has been examined using various theoretical approaches. Mean-field theories work less well near phase transitions because of statistical fluctuations, and this also affects disordered systems [20]. While numerical methods can capture transitions for smaller system sizes in the presence of disorder, they face challenges with finite-size scaling [21].
@@
The Bose glass is both insulating and compressible [22, 23], it has a gapless energy spectrum, and, unlike a superfluid, it has no long-range phase coherence, but instead it develops ‘superfluid puddles’ that are not connected to each other. It was initially believed to exist solely as a ground-state phase [15, 23], however, recent studies suggest that it can also occur at finite temperature [18, 24, 25]. In earlier studies, excitations in a quantum gas following a sudden change (quench) of optical speckle patterns have been observed [26], consistent with the existence of the Bose-glass phase, alongside a reduction of coherence length [27]. Another important question is how disorder affects ergodicity and thermalization. In quantum many-body systems, establishing thermal states locally requires the propagation of correlations, which is restricted to exponentially long timescales when the wavefunctions are localized, either in an insulating or a disordered system. Such phenomena were observed with fermionic and bosonic atoms [28, 29, 30]. Signatures of non-ergodic behavior have also recently been observed in the Bose glass in a two-dimensional optical quasicrystal [31], by probing the adiabaticity of the Bose-glass-to-superfluid transition.
Figure 1: Probing the disordered Bose-Hubbard model. a Illustration of the phase diagram of the 2D disordered Bose-Hubbard model at finite temperature. b Reconstructed occupations of atom clouds in the weakly and strongly interacting regimes with and without disorder. c Schematic to calculate the Edwards Anderson (EA) parameter. For each disorder pattern we record ten fluorescence images and average the occupation on each lattice site to obtain disorder averages, see Eq. (2).
Figure 2: Probing the Bose-glass regime. a Average occupations of ten time-of-flight images at $U/J=15$ and $\Delta/U_{c}=2.8$ (top), $\Delta/U_{c}=0$ (bottom), where $U_{c}$ is the interaction energy at $(U/J)_{c}$, plotted on a logarithmic colorscale for clarity. Blue and black boxes highlight the regions used to determine $n_{\rm max}$ and $n_{\rm min}$, respectively, used to calculate the visibility, $\mathcal{V}$. b Visibility as a function of lattice depth and disorder strength. We have indicated the expected phases SF - Superfluid, MI - Mott insulator, BG - Bose glass. c Edwards-Anderson (EA) parameter (Eq. 2) averaged over the central $5\times 5$ lattice sites (black box in bottom panel of d). The dashed line indicates $\Delta/U=1$, above which the disorder strength is larger than the energy gap in the ideal Mott insulator [23]. d Site-resolved EA parameter, $q_{i}$, at $U/J=55$, using four different disorder realizations with ten in-situ images each, for $\Delta/U_{c}=2.8$ (top) and $\Delta/U_{c}=0$ (bottom). Error bars of $\mathcal{V}$ and $q$ are shown in Figs. S2 and S3 (Supplementary Materials).
@@
In this study, we probe the Bose-glass phase directly using bosonic 87Rb atoms in a quantum-gas microscope, where we project a known, controllable, and repeatable disorder light potential at the scale of individual sites of our square optical lattice [29]. This is in contrast to most other ultra-cold atom experiments on disordered systems, which typically use speckle patterns, quasirandom, or quasicrystalline potentials without precise control over the exact shape of the disorder pattern [32, 33, 34]. Single-site resolved atom detection gives access to the in-situ distribution and particle fluctuations, allowing us to obtain a local measurement of the Edwards-Anderson parameter [35, 36, 37, 38], which qualitatively distinguishes between the Mott insulator and the Bose glass. To distinguish between the superfluid and the Bose glass, we measure the visibility after time of flight to probe long-range phase coherence. By employing Talbot interferometry [39] we find that the coherence length reduces with increasing disorder, consistent with the formation of ‘superfluid puddles’. Finally, we probe adiabaticity when transitioning between the Bose-glass phase and the Mott-insulating or superfluid phases.
Bose glass and Edwards-Anderson parameter
@@
where $J$ is the tunneling strength between neighboring lattice sites, $\hat{a}_{i}^{\dagger}$ and $\hat{a}_{j}$ are the bosonic creation and annihilation operators on neighboring sites $i$ and $j$, respectively, $U$ is the on-site interaction energy, $\hat{n}_{i}=\hat{a}_{i}^{\dagger}\hat{a}_{i}$ is the number operator, $\epsilon_{i}$ is the local energy shift from the harmonic confinement, and $\mu$ is the chemical potential. $\Delta_{i}$ is a local potential that creates random energy offsets on each lattice site, with a uniform box distribution $\Delta_{i}\in[0,\Delta]$, where $\Delta$ is the disorder strength. In the strongly interacting regime, $U\gg J$, a Mott-insulating phase exists for weak disorder $\Delta<U$. When the disorder is larger than the energy gap, $\Delta>E_{g}$, the state becomes gapless [15]. In the weakly interacting regime, the disorder results in localization of atoms and reduces long-range phase coherence. There is no direct transition between the Mott insulator and the superfluid in the presence of disorder for an infinite system [23]. In two dimensions, it is expected that the Bose-glass phase exists between the Mott insulator and superfluid for any non-zero value of $\Delta$ (see phase diagram in Fig. 1a).
@@
In our experiments, we vary $U/J$ by changing the lattice depths $V_{x}$ and $V_{y}$ along the horizontal lattice axes, while keeping the vertical lattice at a depth of $V_{z}=25\,E_{\rm{r}}$ (Methods). We project a known and reproducible repulsive disorder potential generated by a digital mirror device onto a cloud of $\sim 200$ atoms. Using fluorescence imaging, we measure the in-situ site occupations [40]. In the weakly interacting regime, where we expect a superfluid phase in the absence of disorder, we use the interference pattern after a time-of-flight expansion in the vertical lattice to probe the phase coherence of the system [41]. The long-range phase coherence of the superfluid results in peaks in the interference pattern along the horizontal lattice directions. As $\Delta$ is increased, we observe that the visibility, $\mathcal{V}$, of the interference peaks reduces (Fig. 2a) indicating a reduction of global phase coherence. We quantify $\mathcal{V}=\left(n_{\rm max}-n_{\rm min}\right)/\left(n_{\rm max}+n_{\rm min}\right)$, by counting the number of atoms in the peaks ($n_{\rm max}$) and those in regions diagonal from and equidistant to the central peak ($n_{\rm min}$[42]. We probe the superfluid to Bose-glass transition (Fig. 2b) by measuring $\mathcal{V}$ for varying $\Delta$ and $U/J$. For $\Delta=0$, the visibility drops for increasing $U/J$, as a Mott-insulating state starts to form in the center of the cloud. Interestingly, when we increase the disorder strength, we also observe a decrease of $\mathcal{V}$ at $U/J$ below the superfluid-to-Mott-insulator transition, which occurs at $(U/J)_{c}=16.7$ [43, 21]. In the absence of thermal atoms, this reduction in visibility indicates a decreased coherence length, which is consistent with the occurrence of the Bose-glass phase. We compare the visibility measurements to results from Quantum Monte-Carlo simulations (Supplementary Materials) which show the same general trend.
@@
However, the visibility is not a good indicator to identify the transition from a Mott insulator to the Bose glass, as both lack long-range phase coherence. To distinguish between them, we employ the Edwards-Anderson parameter, $q_{i}$, which uses the atom distribution of the in-situ images [37, 38],
@@
Here, the brackets denote the thermal or ensemble average for a single disorder pattern, while the overline represents an average over different disorder patterns. Without disorder, our measurements show $q_{i}\approx 0$ in the center of the cloud (Fig. 2c), limited by thermal excitations at finite temperature. We observe a circular region of non-zero values of $q_{i}$ around the edge of the cloud (Fig. 2d). This is due to the entropy redistribution to the edge of the cloud resulting from the spatially varying chemical potential, and fluctuations in the atom number between repetitions of the experiment. As we increase $\Delta$, the Mott-insulating region decreases until $q_{i}$ starts to rise in the center, indicating that no Mott-insulating component remains. To capture the transition from Mott insulator to Bose glass, we average $q_{i}$ over the central $5\times 5$ lattice sites, $q$, for increasing $\Delta$ (Fig. 2c). When $\Delta>U$, we expect the Bose-glass phase to form across the whole system. We observe $q\approx 0.06$ for our maximum $\Delta$, although the maximum possible value is $q=0.25$. Theoretical studies have shown that at a finite temperature close to ours ($k_{B}T\sim 0.1\,U$), $q$ is significantly reduced from $0.25$ [37, 44], consistent with our experimental observations.
@@
The Edwards-Anderson parameter is unsuitable to distinguish between the superfluid and Bose glass (Fig. 2c), for lattice depths lower than $12\,E_{\rm{r}}$ ($U/J=25.5$), where a superfluid component is present. This disordered superfluid phase has an inhomogeneous density, resulting in a non-zero $q_{i}$, making it indistinguishable from the Bose glass. Therefore, our measurements of the visibility and of the Edwards-Anderson parameter complement each other. Together, they provide evidence of the formation of a Bose glass in the presence of disorder, which we will now probe further.
Talbot interferometry
A distinct feature of the Bose glass is the absence of long-range phase coherence, while maintaining short-range coherence over a small number of sites. To measure the coherence length, we employ an interferometry technique based on the Talbot effect, following an earlier experimental demonstration [39]. It involves briefly switching off the lattice potential in one axis for a duration $t$, before rapidly turning it back on to project the matter wave onto the lattice. At integer multiples of the Talbot time $\tau_{\rm T}=2M\!\,a_{\rm{\ell}}^{2}/h=123\,\upmu$s, where $M$ is the mass of a 87Rb atom, $\,a_{\rm{\ell}}$ is the lattice spacing and $h$ is Planck’s constant, an initially coherent matter wave interferes constructively at the positions of the lattice sites. At non-integer multiples of $\tau_{\rm T}$, the positions of the interfering atoms are no longer aligned with the lattice sites, causing a sudden increase in potential energy when the lattice is turned back on. After a thermalization time, any potential energy manifests itself as an increase in the cloud temperature, resulting in periodic minima of the cloud width at integer multiples of $\tau_{\rm T}$ and periodic maxima at half-integer multiples.

### v2

# Quantum-gas microscopy and Talbot interferometry of the Bose-glass phase

@@ Abstract
Disordered potentials fundamentally affect transport and coherence in quantum systems, giving rise to a Bose-glass phase in interacting bosonic systems - an insulating yet compressible phase lacking long-range coherence. Directly measuring a reduced coherence length of the Bose glass has been a outstanding challenge. We address this by employing Talbot interferometry combined with single-atom-resolved detection in a quantum-gas microscope. Using ultracold bosonic atoms in a two-dimensional lattice with site-resolved, reproducible disorder, we identify the Bose-glass phase through in-situ density distributions and particle-number fluctuations, quantified via the Edwards-Anderson parameter, and through the visibility of interference patterns after time-of-flight. By driving the system across the Bose-glass phase, we further observe signatures of non-ergodic dynamics. Our studies provide a starting point to further explore disordered systems in and out of equilibrium, and are relevant for understanding the dynamics and stability of disordered and glass-like quantum states in solid-state systems.
@@
Disordered potentials govern electron transport and thermalization in real materials [30, 12, 20], and have also been studied in many-body quantum systems with ultracold atoms [51, 58, 47, 3], in both experiment and theory. For non-interacting quantum systems, destructive interference from scattering in the disordered medium leads to Anderson localization [9], a phenomenon that has been observed for example in photonic lattices [57, 32] and in cold-atom systems [50, 44, 29, 28, 67]. In general, the addition of interactions gives rise to a complex set of phenomena including, at low energies, the emergence of the Bose glass [24, 33, 23, 6] and spin glasses [15] in bosonic and fermionic systems. The Bose glass has been examined using various theoretical approaches. Mean-field theories work less well near phase transitions because of statistical fluctuations, and this also affects disordered systems [35]. While numerical methods can capture transitions for smaller system sizes in the presence of disorder, they face challenges with finite-size scaling [60], making the Bose glass an interesting system to study experimentally using a quantum simulator.
@@
The Bose glass is both insulating and compressible [66, 45], it has a gapless energy spectrum, and, unlike a superfluid, it has no long-range phase coherence, but instead it develops ‘superfluid puddles’ that are not connected to each other. It was initially believed to exist solely as a ground-state phase [24, 45], however, recent studies suggest that it can also occur at finite temperature [6, 13, 70]. In earlier studies, excitations in a quantum gas following a sudden change (quench) of optical speckle patterns have been observed [39], consistent with the existence of the Bose-glass phase, alongside a reduction of coherence [41]. Another important question is how disorder affects ergodicity and thermalization. In quantum many-body systems, establishing thermal states locally requires the propagation of correlations, which is restricted to exponentially long timescales when the wavefunctions are localized, either in an insulating or a disordered system. Such phenomena were observed with fermionic and bosonic atoms [56, 18, 48]. Signatures of non-ergodic behavior have also recently been observed in the Bose glass in a two-dimensional optical quasicrystal [68], by probing the adiabaticity of the Bose-glass-to-superfluid transition.
@@
In this study, we probe bosonic atoms in a two dimensional square optical lattice onto which we project a known, controllable, and repeatable disorder light potential at the scale of individual sites [18]. Single-atom resolved detection gives access to the in-situ distribution and particle fluctuations, allowing us to obtain a local measurement of the Edwards-Anderson parameter [22, 40, 62, 2]. We study a regime between the superfluid and Mott-insulating phases in the presence of disorder that shows increased values of the Edwards-Anderson parameter and a reduced visibility in the interference pattern after a time-of-flight expansion. These are complemented by a quantitative measurement of the coherence length using Talbot interferometry. We find that the coherence length reduces with increasing disorder strength, consistent with the formation of ‘superfluid puddles’. Finally, we investigate adiabaticity during transitions between the different regions in the phase diagram, observing non-adiabatic behavior when crossing into the disordered phase, consistent with properties of the Bose-glass phase.
Figure 1: Characterising phases and coherence in the disordered Bose-Hubbard model. a Illustrations of the phases of the disordered Bose-Hubbard model highlighting the degree of coherence between lattice sites. b Schematic to calculate the Edwards-Anderson (EA) parameter. We record ten fluorescence images for a number of disorder patterns at the same disorder strength and average the occupation on each lattice site to obtain disorder averages, see Eq. (2). c We probe the coherence length of the system along one axis using an interferometry technique based on the Talbot effect. Top panel: Time evolution of the sketched one-dimensional initial state after switching off the horizontal lattice. Initially phase coherent regions show (shifted) revivals after (half) integer Talbot times, $\tau_{\rm T}$. Horizontal dashed lines indicate the positions of the lattice sites. Bottom panel: Total energy after recapturing the atoms in the original lattice after a duration $t$. The decay time of the resulting oscillations can be related to the coherence length of the initial state (Appendix).
Figure 2: Probing the disordered Bose-Hubbard model. a Average occupations of ten time-of-flight images at $U/J=15$ and $\Delta/U_{c}=2.8$ (top), $\Delta/U_{c}=0$ (bottom), where $U_{c}$ is the interaction energy at $(U/J)_{c}$, plotted on a logarithmic color scale for clarity. Orange and black boxes highlight the regions used to determine $n_{\rm max}$ and $n_{\rm min}$, respectively, used to calculate the visibility, $\mathcal{V}$. b Visibility as a function of lattice depth and disorder strength. Our analysis identifies superfluid (SF), Mott insulating (MI) and Bose glass (BG) phases. c Edwards-Anderson (EA) parameter (Eq. 2) averaged over the central $5\times 5$ lattice sites (black box in bottom panel of d). The dashed line indicates $\Delta/U=1$, above which the disorder strength is larger than the energy gap in the ideal Mott insulator [45]. d Site-resolved EA parameter, $q_{i}$, at $U/J=55$, using four different disorder realizations with ten in-situ images each, for $\Delta/U_{c}=2.8$ (top) and $\Delta/U_{c}=0$ (bottom). Error bars of $\mathcal{V}$ and $q$ are shown in Figs. S2 and S3 (Supplemental Material).
1 Bose-Hubbard model with disorder
@@
where $J$ is the tunneling strength between neighboring lattice sites, $\hat{a}_{i}^{\dagger}$ and $\hat{a}_{j}$ are the bosonic creation and annihilation operators on neighboring sites $i$ and $j$, respectively, $U$ is the on-site interaction energy, $\hat{n}_{i}=\hat{a}_{i}^{\dagger}\hat{a}_{i}$ is the number operator, $\epsilon_{i}$ is the local energy shift from the harmonic confinement, and $\mu$ is the chemical potential. $\Delta_{i}$ is a local potential that creates random energy offsets on each lattice site, with a uniform box distribution $\Delta_{i}\in[0,\Delta]$, where $\Delta$ is the disorder strength. The model can give rise to a superfluid, a Mott-insulating and a Bose-glass phase (see illustration in Fig. 1a). In the strongly interacting regime, $U\gg J$, a Mott-insulating phase exists for weak disorder $\Delta<U$. When the disorder is larger than the energy gap, $\Delta>E_{g}$, the state becomes gapless [24]. In the weakly interacting regime, the disorder results in localization of atoms and a shorter-range phase coherence compared to a superfluid. In a two-dimensional infinite system it is expected that the Bose-glass phase exists between the Mott insulator and superfluid for any non-zero value of $\Delta$ [45].
2 Experimental details
In our experiments [19], a cloud of ultracold 87Rb atoms is held in a single antinode of a vertical optical lattice (wavelength $\lambda=1064$ nm and depth $V_{z}=25\,E_{\rm{r}}$, where $E_{\rm{r}}=\hbar^{2}/2m\lambda^{2}$ is the recoil energy and $m$ is the mass of a 87Rb atom). After evaporative cooling, a superfluid of $\sim 200$ atoms is prepared in shallow horizontal lattices ($V_{x}=V_{y}\leq 10\,E_{\rm{r}}$). We then vary $U/J$ by changing $V_{x}$ and $V_{y}$ along the horizontal axes. We use fluorescence imaging to obtain the in-situ site occupations of the cloud, which has a diameter of $\sim 16$ sites. Typical temperatures are $k_{B}T\sim 0.1\,U$, measured using radial fits of these occupations in a Mott-insulating state [59].
To create disordered systems, we project a repulsive light potential generated by a blue-detuned laser incident on a digital micromirror device (DMD). At each lattice site, this potential (pictured in Fig. 1b) has a random amplitude, $\Delta_{i}\in[0,\Delta]$, that follows a uniform box distribution (Appendix). Our method of creating disorder patterns is in contrast to most other ultra-cold atom experiments on disordered systems, which typically use speckle patterns, quasirandom, or quasicrystalline potentials [14, 44, 55]. Using the DMD, we can generate not only reproducible disorder patterns, but also maintain phase alignment with the horizontal lattices, which would otherwise drift over time [19].
3 Time-of-flight visibility
In the weakly interacting regime, the presence of disorder should result in a reduced coherence, which we initially measure via the visibility after time-of-flight [10]. This allows us to map out the parameter space for different disorder strengths and lattice depths, within a reasonable experimental runtime. We later complement this with a quantitative measurement of the coherence length via Talbot interferometry (Section 5) for selected parameters.
@@ 3 Time-of-flight visibility
In a time-of-flight measurement, the long-range phase coherence of the superfluid results in peaks in the interference pattern along the horizontal lattice directions. We quantify the visibility of the interference peaks, $\mathcal{V}=\left(n_{\rm max}-n_{\rm min}\right)/\left(n_{\rm max}+n_{\rm min}\right)$, by comparing the number of atoms in the peaks ($n_{\rm max}$, orange boxes in Fig. 2a) to those in regions diagonal from and equidistant to the central peak ($n_{\rm min}$, black boxes) [25]. For $\Delta=0$, the visibility drops for increasing $U/J$ (see bottom row of Fig. 2b), indicating a reduction of global phase coherence as a Mott-insulating state starts to form in the center of the cloud. Interestingly, when we increase $\Delta$, we also observe a decrease in $\mathcal{V}$ at $U/J$ below the superfluid-to-Mott-insulator transition, which occurs at $(U/J)_{c}=16.7$ [16, 60]. In the absence of thermal atoms, this reduction in $\mathcal{V}$ is consistent with the formation of the Bose glass. Our observation of a reduced $\mathcal{V}$ with increasing $\Delta$ cannot be attributed to trivial thermal excitations, as we demonstrate in Section 6. There is no clear transition point in our system due to the underlying harmonic confinement, resulting in a spatially varying chemical potential, and finite-size effects. We also compare the visibility measurements to results from Quantum Monte-Carlo simulations (See Supplemental Material), which show the same general trend.
4 Edwards-Anderson parameter
@@
The visibility is not a reliable observable for distinguishing the Mott insulator from the Bose glass, as both lack long-range phase coherence (rightmost columns of Fig. 2b). Therefore, we employ the Edwards-Anderson parameter, $q_{i}$, which uses the atom distribution of the in-situ images [62, 2],
@@
Here, the brackets denote the thermal or ensemble average for a single disorder pattern, while the overline represents an average over different disorder patterns. Specifically, our data sets are divided into four subsets, one for each disorder pattern (Fig. 1b). For each lattice site $i$, we determine $q_{i}$, which is zero when the occupations resulting from each disorder pattern are equal, and positive when the occupations are different from each other. For consistency, the $\Delta=0$ data were also divided into four subsets, and the small but finite values of $q_{i}$ reflect thermal fluctuations and atom-number variations between these subsets rather than genuine disorder effects.
@@
For $\Delta=0$, our measurements show $q_{i}\approx 0$ in the center of the system (bottom panel of Fig. 2d), limited by thermal excitations at finite temperature, indicative of the Mott insulator. We observe a circular region of non-zero values of $q_{i}$ around the edge of the cloud. This is due to the entropy redistribution to the edge of the cloud resulting from the spatially varying chemical potential, and fluctuations in the total atom number between repetitions of the experiment. As we increase $\Delta$, the Mott-insulating region decreases until $q_{i}$ starts to rise in the center, indicating that no Mott-insulating component remains. To capture the transition from Mott insulator to Bose glass, we average $q_{i}$ over the central $5\times 5$ lattice sites, $q$, for increasing $\Delta$ (Fig. 2c). When $\Delta>U$, the Bose glass is expected to form across the whole system [45]. This predicted behaviour is closely matched by our experimental data, where $q$ increases at disorder strengths $\Delta/U>1$ (black dashed line in Fig. 2c). As we analyze only the center of the system, where the effect of the harmonic confinement is negligible, we see a clearer transition here compared to the global measurement of the visibility (Fig. 2b). We observe $q\approx 0.06$ for our maximum $\Delta$, although the maximum possible value is $q=0.25$. Theoretical studies [62] have shown that at a finite temperature close to ours ($k_{B}T\sim 0.1\,U$), the Edwards-Anderson parameter can still distinguish well between glassy phases and thermal fluctuations. However, its magnitude is significantly reduced from $q=0.25$ [62, 42], consistent with our experimental observations.
@@
The Edwards-Anderson parameter is unsuitable to distinguish between the superfluid and Bose glass, see leftmost columns of Fig. 2c, for lattice depths lower than $12\,E_{\rm{r}}$ ($U/J=25.5$) where a superfluid component is present. This disordered superfluid phase has an inhomogeneous density, resulting in a non-zero $q_{i}$, making it indistinguishable from the Bose glass. Therefore, our measurements of the visibility and of the Edwards-Anderson parameter complement each other. Together, they reveal a region in the phase diagram with properties consistent with the Bose glass. In the measurements that follow, we study the coherence of this region in more detail.
5 Talbot interferometry
[…]